OpenStudy (anonymous):

Suppose the population of a town is 3,400 in 2000. The population decreases at a rate of 2% every 20 years. What will be the projected population in 2040? Round your answer to the nearest whole number.

Mehek (mehek14):

multiply 3400*.02 multiply the answer by 2 then subtract it from 3400

OpenStudy (anonymous):

subtract what from 3400?

Mehek (mehek14):

ok multiply 3400*.02

Mehek (mehek14):

tell me what you get as the answer and then I will tell you what to do next

OpenStudy (anonymous):

3400*.02=68

Mehek (mehek14):

ok good now multiply 68 by 2 because it is 2% for every 20 yrs but they need it for 40 yrs

Mehek (mehek14):

are you still there @onelove26

OpenStudy (anonymous):

okay I got 136

Mehek (mehek14):

ok now subtract 3400 - 136 for the answer

OpenStudy (anonymous):

I got 3,264

OpenStudy (wolf1728):

But it wouldn't be 68 for the second population decrease because it is 2% per 20 years. The difference is small but would affect the answer.

OpenStudy (anonymous):

really? @wolf1728 how would you go about getting the answer then?

OpenStudy (wolf1728):

I'm back --- ^&(&^&%$$# Open Study Okay population in 2020 = 3400 - (3400*.02) = 3,314 Population in 2040 3,314 -(3,314 * .02) = 3,248

Mehek (mehek14):

yea I noticed I messed up on that step

OpenStudy (wolf1728):

So, for 2040 you deduct a population of 66 and not 68

OpenStudy (anonymous):

ohhh

OpenStudy (wolf1728):

Hmm seems I messed up my calculations I think the answer is 3400 - 68 -66 which equals 3,266

OpenStudy (anonymous):

where did you mess that up?

Mehek (mehek14):

at subtraction

OpenStudy (wolf1728):

Originally I typed the answer as 3,248

OpenStudy (wolf1728):

Oh geez I'll have to rethink this population in 2020 = 3,400- (.02 * 3400)= 3,400 - 3,332

OpenStudy (wolf1728):

Population in 2040 = 2020 population minus 2% = 3,332 - (3,332 -.02*3,332) = 3,332 - 66.64 = 3,332 -67 = 3,247

OpenStudy (wolf1728):

Yes, it is a difference of 1 person - but the answer is correct. A differece of 1 is not mush right? Let's suppose you go to the bank and ask them to change a ten dollar bill and they give you 9 one dollar bills? Heck it's only a difference of one right?

OpenStudy (wolf1728):

Thanks for the medal!!